$SPAD/src/input pollevel.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1996.

———-

— * —

\getchunk{license}
-- This file tests variables moving to the correct level within
-- polynomial towers.
)clear all
f(y : UP(x, POLY INT)): UP(x, POLY INT) == x * y
f(x)
)clear all
y := operator 'y
eq1 := D(y x,x,2) - x*y(x)=0
seriesSolve(eq1,y,x=0,[0,a])
eq2 := D(y x,x) - x*cos(y x) - exp(x)
seriesSolve(eq2,y,x=0,y(0) = 0)
)clear all
p : UP(x, POLY INT) := x
u := (z*x)
u * p

———-

References^{[1] nothing}

4